OpenStudy (anonymous):

if f(x)= -3, what is x?

OpenStudy (anonymous):

you have to plug in a number for x

OpenStudy (anonymous):

then you can get multiple answers

OpenStudy (anonymous):

x is any number

OpenStudy (anonymous):

yes

geerky42 (geerky42):

If you graph f(x) = -3, you will get this graph: So no matter the value of x, f(x) is -3. X can be anything.

OpenStudy (freckles):

hmm were you given anything else for the problem? I bet you were given something like Let f(x)=3x+4 if f(x)=-3, what is x? so you replace f(x) with 3x+4 3x+4=-3 <--then you solve for x 3x=-7 x=-7/3 so f(x)=-3 when x=-7/3